Question

You are working in the lab and discover a new factor (X) that appears to play some important role in wound healing. You are provided with the information that each receptor only binds a single X. In order to better characterize the interaction of X with a putative receptor, you obtain dose-response and binding data on X and obtain the following results:

[X]                      Bound X (on 10^7)           Relative Rate of Proliferation

0                             0                                               1

10^-12 M               0.2 pmols                                 1.7

10^-11 M               1.8 pmols                                  4.2

10^-10                   6.0 pmols                                 5.2

10^-9                     10 pmols                                    5.4

10^-8                      12 pmols                                    5.4

These are the data you wish to report:

1. Approximate numbers of X receptor molecules per cell

2. Half-maximal Dose

3. Apparent Kd for the X receptor

4. Any other details that you think might be important

Write a very concise report containing all the information asked and a short paragraph describing why do you think these findings might be interesting.

How would you graph this problem?
Answer #1

1)

Based on the given data,

The receptor and ligands binds in a 1:1 stoichiometry (one ligand per one receptor). Therefore,

10-8 M × 6×1023

= 6×1015 receptors / 107 cells,

So,

6×1015 / 107

X = 6×108 receptors/ cell.

2)

The given data is:

bPnSTeI+ESAcgHYB0AJAOQDoA6QCkA5AOgErpAAB

The dosage response curve is as follows:

2Q==

The Half-maximal Dose is:

y = min + ((max-min)/(1+(X/EC50)-Hill slope))

Where

  • Min = bottom of the curve = 0
  • Max = top of the curve = 12
  • EC50 = Half maximum effective concentration =
  • Hill slope = largest absolute value of the slope of the curve = 3.18

Therefore =

y = 0 + ((12-0)/(1+(6×108/6×1010)-3.18))

= 5.2×10-6

Therefore, the Half maximum does = 5.2×10-6.